Consider a consumer consuming two goods X, and Y, faces the following utility function: U(X,Y)=30X4/5Y1/5 assume further that price X is $5 per unit, price of Y is $10 per unit and income of the consumer is $2000. based on the above information, find the optimum combination of X and Y which maximize the utility
1
Expert's answer
2020-05-13T11:06:31-0400

Utility is maximized at the point where the marginal utility per dollar spent on each good is the same. That is:



"\dfrac{MU_x}{MU_y} = \dfrac{P_x}{P_y}"

The utility function is given as:



"U = 30X^{1\/5}Y^{4\/5}"

The marginal utility for good X is:


"MU_x = 6X^{-4\/5}Y^{4\/5}"

And the marginal utility for good Y is:



"MU_y = 24X^{1\/5}Y^{-1\/5}"

Therefore:



"\dfrac{6X^{-4/5}Y^{4/5}}{ 24X^{1/5}Y^{-1/5}} = \dfrac{5}{10}"

"\dfrac{Y}{X}= 2"

Solving for X and Y each at a time:



"Y = 2X.........(i)"

"X = 0\n.5Y..........(ii)"

The consumer has an income of $2000. Therefore, the budget line is:



"2000 = 5X + 10Y"

Substituting each of the equations above into the budget constraint:



"2000 = 5X + 10(2X)"

"2000 = 25X"

"X ^*= \dfrac{2000}{25} = 80"

"2000 =5(0.5Y) + 10Y"

"2000 =12.5Y"

"Y^* = \dfrac{2000}{12.5} = 160"

Therefore:



"\color{red}{(X^*, Y^*) = (80, 160)}"